can you please help me find the equation of the line parallel to the graph of 3x-y=5 that contains the point (-1,4) and show how you got the answer?
Answer by josmiceli(19441)

Use the point-slope formula.
3x - y = 5
add y to both sides
3x = 5 + y
subtract 5 from both sides and put y on the left side
y = 3x - 5
This is in the form y = mx + b, which is the slope intercept form
m is always the slope
m = 3
Use the point-slope formula for the point (-1, 4)
m = (y - y[1]) / (x - x[1])
3 = (y - 4) / (x + 1)
cross-multiply
3x + 3 = y - 4
add 4 to both sides, put y on the left
y = 3x + 7
check if (-1,4) solves this equation
y = 3*(-1) + 7
y = -3 + 7
y = 4
checks